OpenStudy (anonymous):

Find the remainder when f(x) is divided by (x - k) f(x) = 7x^4 + 12x^3 + 6x^2 - 5x + 16; k = 3

OpenStudy (anonymous):

@gorv @iGreen @phi

OpenStudy (anonymous):

My choices are: 188 946 1704 2512

OpenStudy (gorv):

so we need to find the remainder when divide by x-3

OpenStudy (gorv):

divide it by x-3

OpenStudy (anonymous):

How exactly do I do that?

myininaya (myininaya):

Are you can plug in 3 into f

myininaya (myininaya):

or not are*

myininaya (myininaya):

You can do that by remainder theorem.

OpenStudy (shinalcantara):

using remainder theorem, all you have to do is substitute the value of k to the equation given

OpenStudy (anonymous):

So where it says x I put 3?

OpenStudy (shinalcantara):

yep then evaluate

OpenStudy (shinalcantara):

REMAINDER's THEOREM: \[\frac{ f(x) }{ x-r } = g(x) + \frac{ R }{ x-r }\] where R =f(r)

OpenStudy (anonymous):

I got 946

OpenStudy (shinalcantara):

that's the remainder

OpenStudy (anonymous):

That's all I needed to do for the problem